x+3y=12
-6x+2y=-12

Answer by user_dude2008(1825)

x+3y=12
-6x+2y=-12


6x+18y=72
-6x+2y=-12


20y=60

y=3

x+3y=12
x+3(3)=12
x=12-9
x=3


Ans: x=3 and y=3 ----> (3,3)